OpenStudy (sleepyjess):

Two forces with magnitudes of 150 and 75 pounds act on an object at angles of 30 and 150 respectively. Find the direction and magnitude of the resultant force.

OpenStudy (sleepyjess):

Here is my work so far: y1 = 75*.5*sqrt3 y1 = 64.95 x1 = 75*.5 x1 = 37.5 y2 = 150*.5*sqrt3 y2 = 129.9 x2 = 150*.5 x2 = 75

OpenStudy (sleepyjess):

@Kainui

OpenStudy (kainui):

 See how we can find some nice similarities here? If we cut the right vector into two pieces we really have it as being the other one reflected across the y axis. So their x components cancel and that just leaves their y components doubling. So we just have to add that y component to the vector that's left:  So the new vectors has these components: \[\Large x= 75 \cos( 30 ) \\ \Large y= 3* 75 \sin(30)\] Ok there's a lot going on there, but hopefully that makes sense, if you're not used to playing with vectors maybe this seems weird.